# PICASO: Permutation-Invariant Context Composition with State Space Models

### v1

@@ Abstract
Providing Large Language Models with relevant contextual knowledge at inference time has been shown to greatly improve the quality of their generations. This is often achieved by prepending informative passages of text, or ’contexts’, retrieved from external knowledge bases to their input. However, processing additional contexts online incurs significant computation costs that scale with their length. State Space Models (SSMs) offer a promising solution by allowing a database of contexts to be mapped onto fixed-dimensional states from which to start the generation. A key challenge arises when attempting to leverage information present across multiple contexts, since there is no straightforward way to condition generation on multiple independent states in existing SSMs. To address this, we leverage a simple mathematical relation derived from SSM dynamics to compose multiple states into one that efficiently approximates the effect of concatenating textual contexts. Since the temporal ordering of contexts can often be uninformative, we enforce permutation-invariance by efficiently averaging states obtained via our composition algorithm across all possible context orderings. We evaluate our resulting method on WikiText and MSMARCO in both zero-shot and fine-tuned settings, and show that we can match the strongest performing baseline while enjoying on average 5.4×5.4$\times$5.4 × speedup.
@@ 1 Introduction
Equipping Large Language Models (LLMs) with retrieval capabilities allows them to leverage vast external knowledge bases to greatly improve their generation quality. To achieve this, prepending retrieved sequences to the user prompt is an effective and frequently used approach for incorporating the retrieved information (Ram et al., 2023). This allows the generation to be conditioned on the retrieved data, along with the user prompt.
@@
However, this approach incurs significant computational costs. Not only must the system process the user query and generate an answer, but it must also process the retrieved context, which in real-world settings can amount to thousands of tokens. This problem is exacerbated in Transformer-based models, as the inference cost of generating output tokens scales quadratically with the length of the extended input (see Figure 1).
@@
In contrast, State Space Models (SSMs) offer a more efficient alternative. SSMs encode information from arbitrary-length input sequences into a fixed-size state vector, which can then be conditioned on to generate new tokens without revisiting the original input. This suggests a simple solution: Instead of retrieving from a database containing raw context tokens, we can create a “database of states” containing pre-computed state representations of textual contexts. At inference time, generation starts directly from the retrieved state, simultaneously eliminating the latency from having to process contexts online, and greatly reducing inference time compared to Transformer models (Figure 1).
@@
However, a key challenge arises when conditioning on multiple retrieved contexts. While textual inputs can be simply concatenated, SSM states cannot be composed that easily. To address this, we derive a simple mathematical relation via SSM dynamics to compose states in a manner that exactly equates to the result of concatenation in a single-layer model. Consequently, by simply storing an additional weight matrix along with each context, pre-computed states can be effectively composed at inference time to condition generation on any arbitrary set of contexts.
@@
Since states are computed causally, the order in which contexts are presented affects the state – When there is no natural ordering among retrieved contexts, different order of presentation would yield different states. Consequently, we propose to enforce order-invariance explicitly through averaging states obtained by composing contexts across all possible permutations. While this may appear costly at first glance, we show that the resulting state can be computed exactly in polynomial time (w.r.t. number of contexts) using Dynamic Programming, and this can be further reduced to linear time by accepting slight approximations. This greatly benefits Retrieval-Augmented Generation (RAG) tasks, where our results show a $10\%$ improvement over the best order-dependent state composition method when order-invariance is incorporated into the conditioned state.
@@
To summarize, we introduce a method for efficiently retrieving and composing multiple pre-computed states at inference time to condition the generation of high-quality outputs, which we term PICASO (Permutation-Invariant Compositional Aggregation of States as Observations). Our experiments show that PICASO achieves 91% of the performance gain from combining the raw tokens of multiple contexts, while offering a 5.4×5.4$\times$5.4 × speed-up over concatenation.
@@ 2 Related Work
More recently, many of these ideas have been rediscovered and implemented on modern parallel hardware as basic building blocks for Foundation Models. Gu & Dao (2023) proposed Mamba, a selective (input-dependent) SSM based on LIV systems, demonstrating comparable performance to Transformers (Vaswani, 2017) on language modeling while enjoying faster inference. Mamba-2 (Dao & Gu, 2024) further improved computational time by implementing SSM layers with structured matrix multiplications to better leverage modern Tensor Cores. Although pure SSM models can compete with Transformer blocks on many NLP tasks, they lag behind on tasks that require strong recall capabilities. To balance inference efficiency and model capabilities, Hybrid models combining Attention and SSM blocks have been proposed. Lieber et al. (2024) combined SSM blocks along with global-attention blocks to create a hybrid architecture with Mixture-of-Expert layers for training larger models. To further improve long context ability and efficiency, Ren et al. (2024) leveraged sliding window attention, while Zancato et al. (2024) developed a generic architecture that generalizes Hybrid models, allowing effective access to both eidetic and fading memory.
@@ 3.1 Preliminaries:
A linear input-dependent discrete-time model has the form
@@ 6.1 Implementation Details
We run our main experiments on the largest available SSM on Huggingface - Mamba-2 2.7B (Dao & Gu, 2024). We evaluate our method on two large-scale datasets - WikiText-V2 (Merity et al., 2016) and MSMARCO (Nguyen et al., 2016). We use the training splits as our fine-tuning data, and the testing/validation splits respectively for evaluation. To pre-process WikiText-V2 for our use case, we split each passage in the dataset into two equal context “chunks”, with the goal of predicting the second (continuation) from the first (query). The retrieval database comprises all remaining chunks, from which we retrieve via an external sentence embedding model, All-MiniLM-L6-v2111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. In most experiments, we retrieve up to 10 chunks, since improvements appears to saturate beyond that, and loss from concatenation blows up as a result of exceeding training context length (Figure 6, Appendix). We pre-process MSMARCO by filtering only entries with well-formed answers and discarding those without relevant passages.
@@
For fine-tuning experiments using BPTC and BP2C, we base our implementation on the official HuggingFace  [^3] https://github.com/huggingface/transformers trainer with default hyperparameters, and retrieve the $k$ most relevant context chunks for each query sample for composition. For WikiText, we select $k\in\{0,\ldots,10\}$ uniformly at random for each batch. For MSMARCO, we use all the available passages (both relevant and irrelevant) associated with each training example. For both datasets, we fine-tune for only 1 epoch. In all fine-tuning experiments, we ensure the training set (both the examples and the chunk database) are disjoint from the validation set to ensure fair evaluation.
@@ 6.2 Comparison Models
Concatenation (Ram et al., 2023): We concatenate individual context chunks based on a specific ordering. For WikiText-V2 experiments, we consider the “best-case ordering” as determined by the sentence embedding model where more relevant contexts are closer to the query (at the end). We initialize the model with the state of the earliest context chunk in the concatenation, which we assume to be available via pre-processing, and recompute the composed state from only the remaining ones.
@@ 6.3.1 Zero-shot performance
We demonstrate in Figure 3 that applying PICASO-R in a zero-shot manner on WikiText-V2 greatly improves performance over the baseline by an average of $10.1\%$ across 1-10 retrieved context chunks. This greatly improves over Soup ($8.5\%$) and CASO ($9.2\%$). Compared to concatenation ($11.1\%$), PICASO-R performs slightly worse but benefits from magnitudes improvement in processing time on an average of 5.4×5.4$\times$5.4 ×. In this task, PICASO-R achieves almost exactly the same performance as PICASO-S, but with a much faster composition time. As a sanity check for motivation for our method, we show that PIConcat achieves the best performance ($12.0\%$) overall, but at the cost of significantly greater computational time despite our batched-inference implementation.
@@ 6.3.2 Backpropagation Through and To Composition
Table 1: All models in this table are evaluated on the MSMARCO validation set. We evaluate performance of models fine-tuned via BPTC/BP2C on both the WikiText and MSMARCO training sets. Rows 2-3 show that fine-tuning models to compose WikiText context chunks does not harm performance when evaluated on composing context chunks from MSMARCO. When composing chunks from distributions similar to those encountered during training (Rows 4-5), PICASO matches the performance of concatenation while being magnitudes faster.
@@ 7 Limitations and Discussion
We have proposed a method, PICASO, that enables efficient retrieval and composition of contexts by pre-processing their individual states. Without any training, our approach can handle the composition of information contained in up to 10 context chunks in a manner that is order-invariant. PICASO notably requires zero online model processing time, since generation can begin directly from the composed states. When models are further fine-tuned with our proposed learning objective, states composed using PICASO perform comparably to those produced from the concatenation of context tokens, while offering on average a 5.4×5.4$\times$5.4 × faster composition time.
@@ B.2 Scaling beyond effective context length
Figure 6: Concatenation scales poorly with total size of retrieved contexts beyond training context length. PICASO yields greater stability even composing up to 50 context chunks retrieved from WikiText.
In Figure 6, we show that as the total length of retrieved contexts scale beyond a certain threshold (effective context size of the model), the loss from concatenation blows up and rapidly increases beyond the no-retrieval baseline. On the other hand, performance of PICASO remains stronger than that of the baseline when composing 50 context chunks.
@@ B.6 Evaluation on Multiple Choice Tasks
Table 3: Evaluation on OpenbookQA dataset, augmented with retrieved passages from WikiText. We use normalized accuracy as our evaluation metric, and report the time taken to compose retrieved passages. Numbers shown in the table are averaged across retrieving between 1 to 10 full contexts from WikiText (as opposed to half context chunks in our main paper experiments).
@@ B.7 Context Statistics
In Figure 9, we plot the distribution over the lengths (tokens and characters) of retrieved context chunks used in the main paper WikiText retrieval dataset.
Figure 9: Histogram of the lengths, in terms of (Left) characters and (Right) tokens, of database context chunks used in the main paper WikiText experiments.
@@ Appendix C Data Attribution
The first method, which we term ”Leave-one-in”, is to prepend each candidate context to the question, and evaluate the loss on the answer. Equivalently, $\operatorname*{arg\,min}_{i}L_{CE}(f_{\theta}(\bm{u}_{q},x(\bm{u}_{i})),\bm{u}_{a})$, where we abuse notation to denote loss on the sequence (instead of token) $\bm{u}_{a}$.
The second method, which we term ”Leave-one-out”, is to compare the marginal increase in loss of the answer when removing each candidate from the composition of all of them. Equivalently, $\operatorname*{arg\,max}_{i}\{L_{CE}(f_{\theta}(\bm{u}_{q},\hat{x}(S_{-i})),\bm{u}_{a})-L_{CE}(f_{\theta}(\bm{u}_{q},\hat{x}(S)),\bm{u}_{a})\}$, where $\hat{x}(S_{-i})$ denotes a state composed from all contexts in $S$ other than $\bm{u}_{i}$.
@@ Appendix D Concatenation for SSMs: Connection to jump-linear systems
Consider a collection of context segments retrieved based on relevance to a query, and sorted randomly as context to the query. While these segments, or ‘chunks’ share information, they are independent given the query, and their order is accidental and uninformative.
We are interested in a model that can efficiently process inputs in this format and extract all shared information from the input. Attention-based models are a natural choice because of the permutation-invariance of attention mechanisms (ignoring positional encoding), but they would have to process the entire input (all chunks) with quadratic inference cost. On the other hand, SSMs have linear cost, but they are ill-fit to process this kind of input because of the context switches, which make the Markov assumption implicit in the state representation invalid.
@@
Since modern state space models are input-dependent, they automatically fit the latter class of models and can handle switches without modifications. However, what they cannot handle is the fact that the order of the chunks is uninformative. As a result, presenting the same chunks in different order would yield different states. Accordingly, our goal is to enable SSMs to learn from segments up to permutations, so we can accommodate sequences where the ordering within chunks is informative and respected, while the ordering of chunks is uninformative and factored out.

### v2

@@ Abstract
Providing Large Language Models with relevant contextual knowledge at inference time has been shown to greatly improve the quality of their generations. This is often achieved by prepending informative passages of text, or ‘contexts’, retrieved from external knowledge bases to their input. However, processing additional contexts online incurs significant computation costs that scale with their length. State Space Models (SSMs) offer a promising solution by allowing a database of contexts to be mapped onto fixed-dimensional states from which to start the generation. A key challenge arises when attempting to leverage information present across multiple contexts, since there is no straightforward way to condition generation on multiple independent states in existing SSMs. To address this, we leverage a simple mathematical relation derived from SSM dynamics to compose multiple states into one that efficiently approximates the effect of concatenating raw context tokens. Since the temporal ordering of contexts can often be uninformative, we enforce permutation-invariance by efficiently averaging states obtained via our composition algorithm across all possible context orderings. We evaluate our resulting method on WikiText and MSMARCO in both zero-shot and fine-tuned settings, and show that we can match the strongest performing baseline while enjoying on average 5.4×5.4$\times$5.4 × speedup.
@@ 1 Introduction
Incorporating new information in deep learning models has traditionally been a costly process, often requiring re-training or fine-tuning their weights on new data. Fortunately, Large Language Models (LLMs) provide a compelling alternative: These models can ‘learn’ to leverage new contextual information in real-time by simply prepending them as inputs, without having to modify their weights (Ram et al., 2023). This has motivated a powerful application known as Retrieval-Augmented Generation (RAG), where LLMs are deployed with the ability to retrieve and incorporate relevant sources of information, or ‘contexts’, from vast external knowledge bases when queried by users at inference time.
@@
Despite being faster than the naive fine-tuning of model weights, this approach still incurs significant computational costs. Not only must the system process the user query and generate an answer, but it must also process the retrieved context, which in real-world settings can amount to thousands of tokens. This problem is exacerbated in Transformer-based models, as the inference cost of generating output tokens scales quadratically with the length of the extended input (see Figure 1).
@@
In contrast, State Space Models (SSMs) offer a more efficient alternative. SSMs encode information from arbitrary-length input sequences into a fixed-size state vector, which can then be conditioned on to generate new tokens without revisiting the original input. This suggests a simple method to reduce the cost of incorporating new contexts at inference-time: Instead of retrieving from a database containing raw context tokens, we can create a “database of states” containing pre-computed state representations of contexts. At inference time, generation starts directly from the retrieved state, simultaneously eliminating the latency from having to process context tokens online, and greatly reducing inference time compared to Transformer models (Figure 1).
@@
However, a key challenge arises when conditioning on multiple retrieved states. While input tokens can be simply concatenated and fed into an LLM or SSM, existing models are trained to generate outputs conditioned only on a single SSM state. To address this, we derive a simple mathematical relation via SSM dynamics to compose multiple states into one, in a manner that exactly equates to the result of concatenation in a single-layer model. Consequently, by simply storing an additional weight matrix along with each context, pre-computed states can be effectively composed at inference time to condition generation on any arbitrary set of contexts.
@@
Since states are computed causally, the order in which contexts are presented affects the state – when there is no natural ordering among retrieved contexts, different order of presentation would yield different states. Consequently, we propose to enforce order-invariance explicitly through averaging states obtained by composing contexts across all possible permutations. While this may appear costly at first glance, we show that the resulting state can be computed exactly in polynomial time in the number of context segments using Dynamic Programming, and this can be further reduced to linear time by accepting slight approximations. This greatly benefits Retrieval-Augmented Generation tasks, where our results show a $10\%$ improvement over the best order-dependent state composition method when order-invariance is incorporated into the conditioned state.
@@
To outline our main contributions, we introduce a method for efficiently retrieving and composing multiple pre-computed states at inference time to condition the generation of high-quality outputs, which we term PICASO (Permutation-Invariant Compositional Aggregation of States as Observations). Our experiments show that PICASO achieves 91% of the performance gain from combining the raw tokens of multiple contexts, while offering a 5.4×5.4$\times$5.4 × speed-up over concatenation.
@@ 2 Related Work
More recently, many of these ideas have been rediscovered and implemented on modern parallel hardware as basic building blocks for Foundation Models. Gu & Dao (2023) proposed Mamba, an input-dependent linear SSM (termed ‘selective’) based on LIV systems, that achieves comparable performance to Transformers (Vaswani, 2017) on language modeling while enjoying faster inference. Mamba-2 (Dao & Gu, 2024) further improved computational time by implementing SSM layers with structured matrix multiplications to better leverage modern Tensor Cores. Although pure SSM models can compete with Transformer blocks on many NLP tasks, they lag behind on tasks that require strong recall capabilities. To balance inference efficiency and model capabilities, Hybrid models combining Attention and SSM blocks have been proposed. Lieber et al. (2024) combined SSM blocks along with global-attention blocks to create a hybrid architecture with Mixture-of-Expert layers for training larger models. To further improve long context ability and efficiency, Ren et al. (2024) leveraged sliding window attention, while Zancato et al. (2024) developed a general family of architecture that include Transformers, SSMs and their hybrid combinations, leveraging both verbatim and fading memory, in both long- and short-term.
@@ 3.1 Preliminaries:
A linear input-dependent discrete-time state-space model has the form
@@ 6.1 Implementation Details
We run our main experiments on the largest available SSM on Huggingface - Mamba-2 2.7B (Dao & Gu, 2024). We evaluate our method on two large-scale datasets - WikiText-V2 (Merity et al., 2016) and MSMARCO (Nguyen et al., 2016). We use the training splits as our fine-tuning data, and the testing/validation splits respectively for evaluation. To pre-process WikiText-V2 for our use case, we split each passage in the dataset into two equal context “segments”, with the goal of predicting the second (continuation) from the first (query). The retrieval database comprises all remaining segments, from which we retrieve via an external sentence embedding model, All-MiniLM-L6-v2111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. In most experiments, we retrieve up to 10 segments, since improvements appears to saturate beyond that, and loss from concatenation blows up as a result of exceeding training context length (Figure 6, Appendix). We pre-process MSMARCO by filtering only entries with well-formed answers and discarding those without relevant passages.
@@
For fine-tuning experiments using BPTC and BP2C, we base our implementation on the official HuggingFace  [^3] https://github.com/huggingface/transformers trainer with default hyperparameters, and retrieve the $k$ most relevant context segments for each query sample for composition. For WikiText, we select $k\in\{0,\ldots,10\}$ uniformly at random for each batch. For MSMARCO, we use all the available passages (both relevant and irrelevant) associated with each training example. For both datasets, we fine-tune for only 1 epoch. In all fine-tuning experiments, we ensure the training set (both the examples and the context database) are disjoint from the validation set to ensure fair evaluation.
@@ 6.2 Comparison Models
Concatenation (Ram et al., 2023): We concatenate individual context segments based on a specific ordering. For WikiText-V2 experiments, we consider the “best-case ordering” as determined by the sentence embedding model where more relevant contexts are closer to the query (at the end). We initialize the model with the state of the earliest context segment in the concatenation, which we assume to be available via pre-processing, and recompute the composed state from only the remaining ones.
@@ 6.3.1 Zero-shot performance
We demonstrate in Figure 3 that applying PICASO-R in a zero-shot manner on WikiText-V2 greatly improves performance over the baseline by an average of $10.1\%$ across 1-10 retrieved context segments. This greatly improves over Soup ($8.5\%$) and CASO ($9.2\%$). Compared to concatenation ($11.1\%$), PICASO-R performs slightly worse but benefits from magnitudes improvement in processing time on an average of 5.4×5.4$\times$5.4 ×. In this task, PICASO-R achieves almost exactly the same performance as PICASO-S, but with a much faster composition time. As a sanity check for motivation for our method, we show that PIConcat achieves the best performance ($12.0\%$) overall, but at the cost of significantly greater computational time despite our batched-inference implementation.
@@ 6.3.2 Backpropagation Through and To Composition
Table 1: All models in this table are evaluated on the MSMARCO validation set. We evaluate performance of models fine-tuned via BPTC/BP2C on both the WikiText and MSMARCO training sets. Rows 2-3 show that fine-tuning models to compose WikiText context segments does not harm performance when evaluated on composing context segments from MSMARCO. When composing segments from distributions similar to those encountered during training (Rows 4-5), PICASO matches the performance of concatenation while being magnitudes faster.
@@ 7 Limitations and Discussion
We have proposed a method, PICASO, that enables efficient retrieval and composition of contexts by pre-processing their individual states. Without any training, our approach can handle the composition of information contained in up to 10 context segments in a manner that is order-invariant. PICASO notably requires zero online model processing time, since generation can begin directly from the composed states. When models are further fine-tuned with our proposed learning objective, states composed using PICASO perform comparably to those produced from the concatenation of context tokens, while offering on average a 5.4×5.4$\times$5.4 × faster composition time.
@@ B.2 Scaling beyond effective context length
Figure 6: Concatenation scales poorly with total size of retrieved contexts beyond training context length. PICASO yields greater stability even composing up to 50 context segments retrieved from WikiText.
In Figure 6, we show that as the total length of retrieved contexts scale beyond a certain threshold (effective context size of the model), the loss from concatenation blows up and rapidly increases beyond the no-retrieval baseline. On the other hand, performance of PICASO remains stronger than that of the baseline when composing 50 context segments.
@@ B.6 Evaluation on Multiple Choice Tasks
Table 3: Evaluation on OpenbookQA dataset, augmented with retrieved passages from WikiText. We use normalized accuracy as our evaluation metric, and report the time taken to compose retrieved passages. Numbers shown in the table are averaged across retrieving between 1 to 10 full contexts from WikiText (as opposed to half context segments in our main paper experiments).
@@ B.7 Context Statistics
In Figure 9, we plot the distribution over the lengths (tokens and characters) of retrieved context segments used in the main paper WikiText retrieval dataset.
Figure 9: Histogram of the lengths, in terms of (Left) characters and (Right) tokens, of database context segments used in the main paper WikiText experiments.
@@ Appendix C Data Attribution
The first method, which we term “Leave-one-in”, is to prepend each candidate context to the question, and evaluate the loss on the answer. Equivalently, $\operatorname*{arg\,min}_{i}L_{CE}(f_{\theta}(\bm{u}_{q},x(\bm{u}_{i})),\bm{u}_{a})$, where we abuse notation to denote loss on the sequence (instead of token) $\bm{u}_{a}$.
The second method, which we term “Leave-one-out”, is to compare the marginal increase in loss of the answer when removing each candidate from the composition of all of them. Equivalently, $\operatorname*{arg\,max}_{i}\{L_{CE}(f_{\theta}(\bm{u}_{q},\hat{x}(S_{-i})),\bm{u}_{a})-L_{CE}(f_{\theta}(\bm{u}_{q},\hat{x}(S)),\bm{u}_{a})\}$, where $\hat{x}(S_{-i})$ denotes a state composed from all contexts in $S$ other than $\bm{u}_{i}$.
@@ Appendix D Concatenation for SSMs: Connection to jump-linear systems
Consider a collection of context segments retrieved based on relevance to a query, and sorted randomly as context to the query. While these segments share information, they are independent given the query, and their order is accidental and uninformative.
We are interested in a model that can efficiently process inputs in this format and extract all shared information from the input. Attention-based models are a natural choice because of the permutation-invariance of attention mechanisms (ignoring positional encoding), but they would have to process the entire input (all segments) with quadratic inference cost. On the other hand, SSMs have linear cost, but they are ill-fit to process this kind of input because of the context switches, which make the Markov assumption implicit in the state representation invalid.
@@
Since modern state space models are input-dependent, they automatically fit the latter class of models and can handle switches without modifications. However, what they cannot handle is the fact that the order of the segments is uninformative. As a result, presenting the same segments in different order would yield different states. Accordingly, our goal is to enable SSMs to learn from segments up to permutations, so we can accommodate sequences where the ordering within segments is informative and respected, while the ordering of segments is uninformative and factored out.